OpenStudy (anonymous):

Hi everyone! Can someone help me solve a Laplace transform? f(t)=e^(t)sinh(t)... Pretty confused and I can't seem to find an example anywhere. Thanks! :o)

zepdrix (zepdrix):

:D

OpenStudy (anonymous):

sorry, I wrote the wrong question initially but it is now correct and still need help.

zepdrix (zepdrix):

Remember how to rewrite sinh(t) in terms of exponentials? :)

OpenStudy (anonymous):

is that the only way to solve this?

zepdrix (zepdrix):

I don't see why you want to do any other method :D hmm

OpenStudy (anonymous):

well what I mean is this...gimme a sec to explain and maybe you can correct my thinking :o)...

OpenStudy (michele_laino):

by definition of Laplace transform, we can write: \[\Large F\left( s \right) = \int_0^\infty {{e^t}\sinh \left( t \right){e^{ - st}}} dt\]

zepdrix (zepdrix):

Ahh it got cut off >.< But something like that, ya? :) Just taking Laplace of two exponentials.

zepdrix (zepdrix):

\[\large\rm \mathcal{L}\{ e^t \sinh(t)\}=\mathcal{L}\{ e^t \cdot\frac{1}{2}(e^t-e^{-t})\}=\frac{1}{2}\left(\mathcal{L}\{e^{2t}\}-\mathcal L\{e^{0}\}\right)\]

OpenStudy (michele_laino):

now, using the suggestion of @zepdrix we can make this sbstitution: \[\Large \sinh t = \frac{{{e^t} - {e^{ - t}}}}{2}\]

OpenStudy (anonymous):

when I first tried solving the problem, I first looked at e^t and recognized it of the form L{e^at}=1/s-a Then I looked at L{sinht} and recognized it as k=1 so therefore K/s^2-K^2 which would equal 1/s^2+1 so then I thought I could multiply them together>>> (1/s-a) times (1/s^2+1) to get 1/(s-a)(s^2-1) but that isn't the answer... first of all, can you see why I would want to solve it like this? second, can you tell me exactly why this is the wrong method? :o)

zepdrix (zepdrix):

Hmm are you confusing sin(t) with sinh(t) maybe? :)

OpenStudy (anonymous):

no...i'm not confusing them L{sinkt}=k/s^2+k^2 whereas L{sinhkt}=k/s^2-k^2 all I was trying to do was use simple definitions of some common Laplace transforms

zepdrix (zepdrix):

Oh yes your formula looks good hehe :)

OpenStudy (anonymous):

can you explain why I can't do the problem this way?

OpenStudy (anonymous):

the only thing I can think of is that you can't take L{ab} and change it to L{a}timesL{b} is that correct?

zepdrix (zepdrix):

Well I mean.. you can't do that in general with integration. Example:\[\Large\rm \int\limits \frac{1}{2}x~dx=x^2,\qquad\qquad \int\limits \cos(x)dx=\sin(x)\]But,\[\Large\rm \int\limits \frac{1}{2}x \cos(x)~dx\ne x^2 \sin(x)\]

zepdrix (zepdrix):

Yah, the last thing you said sounds correct :)

zepdrix (zepdrix):

Laplace Transform is a `linear operator` so it satisfies these two properties: L{a+b} = L{a} + L{b} L{ca} = cL{a} (Where c is a constant). Wish that multiplication worked that way though! hehe would be nice

OpenStudy (anonymous):

haha...okay...bleh :o/ then I just have some simple rules mixed up...so then if I can't solve it remembering that sinht=(that e^2 thingy) then I can do the long integral form that Michele posted above correct?

zepdrix (zepdrix):

Yah you might be able to do that integral by parts, remembering that your derivative of sinh(t) is cosh(t). And you'd have to do parts like 2 times.. and then that weird algebra step thing... if you remember back to this integral: \(\Large\rm \int e^x\sin(x)dx\). Kinda how you dealt with that guy.

zepdrix (zepdrix):

But converting to exponentials seems way better XD hehe

OpenStudy (michele_laino):

hint: you vcan not multiply the single laplace transforms, since what you get is the laplace transform of the convolution product of the functions, which the single laplace transforms are referring to

OpenStudy (michele_laino):

can*

OpenStudy (anonymous):

sorry...the website is messing up

OpenStudy (anonymous):

okay yes...I would eventually have to add or subtract the laplace transforms and then factor it out and all that ghastly algebra hoping I don't flip a sign ?

OpenStudy (anonymous):

okay...so the key here is remembering that the Laplace transform itself is a linear operation, so no multiplication! that's a no-no correct?

OpenStudy (anonymous):

<8-/ ?

zepdrix (zepdrix):

ya linear operation gives us those `two properties` but unfortunately not that other one you were hoping for :D

OpenStudy (anonymous):

are you guys still here?

zepdrix (zepdrix):

These are really handy to remember though, you should tryyyyy!\[\Large\rm \sin(x)=\frac{1}{2i}\left(e^{ix}-e^{-ix}\right),\qquad \sinh(x)=\frac{1}{2}\left(e^{x}-e^{-x}\right)\]\[\Large\rm \cos(x)=\frac{1}{2}\left(e^{ix}+e^{-ix}\right),\qquad \cosh(x)=\frac{1}{2}\left(e^{x}+e^{-x}\right)\]Just tryyyy >.< They match up so nicely!

OpenStudy (michele_laino):

that's right! @SinginDaCalc2Blues

OpenStudy (anonymous):

sorry...I think this website is really messing up...I couldn't see anything you posted until now...gimme just a sec to review :o)

zepdrix (zepdrix):

ya it tweaked out for a second there >.<

OpenStudy (anonymous):

hey zepdrix, are you missing an"i" next to the 2 in cost above?

zepdrix (zepdrix):

you mean the cos(x) identity? :o No. no i in the denominator if that's what you're asking.

OpenStudy (anonymous):

so the sinx has an "i" in the denominator but not the cosx?

zepdrix (zepdrix):

ya :D Do you remember ummm... Euler's Formula thingy? e^(ix) = cos x + i sin (x) Sine has the imaginary unit on it. So when you do a little bit of fancy math... . . . the sine identity is the only one that ends up with the i.

OpenStudy (anonymous):

I never commited that euler thingy to memory, I always used the alternate e^rx(cos bla bla stuff)

zepdrix (zepdrix):

ya it's not super important until you take complex analysis. up until that point, it's really just a "neato" formula :D

OpenStudy (anonymous):

now I can't even remember what the bla bla stuff is now that I think about it...uhg! e^rx[(C1cos(betax) + C2sin(betax)] is that right?

OpenStudy (anonymous):

that was for imaginary roots

zepdrix (zepdrix):

Mmm I'm not sure what that is. Oh oh yes, when you end up with complex roots to a ODE :) Ya ya looks familiar.\[\Large\rm y=c_1 e^{ax}\sin(bx)+c_2 e^{ax}\cos(bx)\]Where your roots to your characteristic were something like \(\Large\rm r=a+bi\)

zepdrix (zepdrix):

Something like that, ya >.<

OpenStudy (anonymous):

yes that's it...i misspoke...it is complex root, not imaginary roots

OpenStudy (anonymous):

so unless I want to do the long integral, I need to remember the sinht and cosht then

OpenStudy (anonymous):

omg...just noticed...3000 medal for your label! wow i'm happy with my little 50 medals for human calculator :o)

zepdrix (zepdrix):

oh haha XD ya too much free time i guess lolol

zepdrix (zepdrix):

ya just remmeber them :UUU

zepdrix (zepdrix):

i dont see any way around it :3

OpenStudy (anonymous):

are you still a student also? how do you have that much time to help people?

OpenStudy (anonymous):

very impressive!

zepdrix (zepdrix):

that's a good question :) i dunno.. i just love teaching i guess.. always find time for it. i do tutoring for a part time job, and i come home and tutor for free on the internet, do what you love i guess ;)

OpenStudy (anonymous):

that's really neat! thank you for all you help and don't say thanks! :o)

zepdrix (zepdrix):

XD

OpenStudy (anonymous):

ok so back to my problem... change e^tsinht into e^t(e^x-e^-x/2)...then combine the base "e's"

zepdrix (zepdrix):

ya e^t((e^t-e^-t)/2)

zepdrix (zepdrix):

Ya exponent rule :U

OpenStudy (anonymous):

okay.... still working on it...

OpenStudy (anonymous):

so I could break it down like this I suppose... e^t(e^x - e^-x)(1/2) then (1/2)(e^xt) - (1/2)(e^-xt) then (1/2)L{e^xt} - (1/2)L{e^-xt} a=x....................a=-x (1/2)(1/s-x) - (1/2)(1/s+x) how does this look so far?

zepdrix (zepdrix):

Mmm no no no. Now you're just being silly :o) You were dealing with e^t sinh(t) NOT WITH e^t sinh(x)

OpenStudy (anonymous):

okay well replace the x with t then

OpenStudy (anonymous):

wait no

zepdrix (zepdrix):

And make sure you remember your exponent rules correctly! :D\[\Large\rm e^t\cdot e^x\ne e^{xt}\]Even though that's not what we were looking for anyway :d

OpenStudy (anonymous):

ugh...just a sec...lemme redo

zepdrix (zepdrix):

haha k XD

OpenStudy (anonymous):

okay...hopefully I just found my misplaced remaining brain cell... e^t(e^x - e^-x)(1/2) then (1/2)(e^(t+x)) - (1/2)(e^(t-x)) then (1/2)L{e^(t+x)} - (1/2)L{e^(t-x)} a=(t+x)....................a=(t-x) (1/2)(1/s-(t+x)) - (1/2)(1/s+(t-x)) how does this look so far?

OpenStudy (anonymous):

wait no

zepdrix (zepdrix):

AHH I wish you would get those stupid x's out :U

zepdrix (zepdrix):

they're all t's! x does exist! >.<

OpenStudy (anonymous):

i needed to factor out a t right?

zepdrix (zepdrix):

I dunno, you're being a silly billy. We need to slow this down maybe..

OpenStudy (anonymous):

what do you mean the x's don't exist? I thought sinht was e^x-e^-x/2 ? should it be instead e^t-e^-t/2 ?

zepdrix (zepdrix):

Yes silly! :U\[\Large\rm \sinh(x)=\frac{1}{2}(e^x-e^{-x}),\qquad\sinh(t)=\frac{1}{2}(e^t-e^{-t})\]

OpenStudy (anonymous):

omg! of course

OpenStudy (anonymous):

okay...and just to make sure, to get it in the form of L{e^at} i need to factor out the t's right?

OpenStudy (anonymous):

well maybe since I got rid of the x's, if I back the problem up, it may become more simple...one sec

OpenStudy (anonymous):

i'm not gonna look at your work...I need to figure it out...can you erase that please?

OpenStudy (anonymous):

okay...one more try here...gimme a sec

zepdrix (zepdrix):

SinginDaODE Blueeeees

OpenStudy (anonymous):

shaaa

zepdrix (zepdrix):

Figure it out lady blues? :D

OpenStudy (anonymous):

e^t(e^t - e^-t)(1/2) then (1/2)(e^(t+t)) - (1/2)(e^(t-t)) then (1/2)L{e^(2t)} - (1/2)L{e^(0)} then (1/2)L{e^(2t)} - (1/2)L{1} a=(2) (1/2)(1/s-2) - (1/2)(1/s) then I suppose you could write the answer in many different ways but... (1/2) [ (1/s-2) -(1/s) ]

zepdrix (zepdrix):

yayyy \c:/ good job!

OpenStudy (anonymous):

OMG...whew :o)

zepdrix (zepdrix):

ok im goin to bed -_- its too late

OpenStudy (anonymous):

thanks @zepdrix ! :o)

OpenStudy (anonymous):

3001 medals for u! :o)